OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (1, 3) m = 3

OpenStudy (studygurl14):

I'm pretty sure I helped you with a VERY similar one...Try this one by yourself. What do you do first? y = mx + b m = 3

OpenStudy (anonymous):

first plug in the point in the correct spot

OpenStudy (anonymous):

3=3(1)+b 3=3+b -3 -3 0=b